# Does differentially private synthetic data lead to synthetic discoveries?

### v1

@@ Abstract
Objectives: The aim of this study is to evaluate the Mann-Whitney U test on DP-synthetic biomedical data in terms of Type I and Type II errors, in order to establish whether statistical hypothesis testing performed on privacy preserving synthetic data is likely to lead to loss of test’s validity or decreased power.
Methods: We evaluate the Mann-Whitney U test on DP-synthetic data generated from real-world data, including a prostate cancer dataset (n=500) and a cardiovascular dataset (n=70 000), as well as on data drawn from two Gaussian distributions. Five different DP-synthetic data generation methods are evaluated, including two basic DP histogram release methods and MWEM, Private-PGM, and DP GAN algorithms.
Conclusion: Most of the tested DP-synthetic data generation methods showed inflated Type I error, especially at privacy budget levels of $\epsilon\leq 1$. This result calls for caution when releasing and analyzing DP-synthetic data: low p-values may be obtained in statistical tests simply as a byproduct of the noise added to protect privacy. A DP smoothed histogram-based synthetic data generation method was shown to produce valid Type I error for all privacy levels tested but required a large original dataset size and a modest privacy budget ($\epsilon\geq 5$) in order to have reasonable Type II error levels.
@@ 1 Introduction
Hypothesis tests are essential statistical tools in analyzing biomedical data. Mann-Whitney (MW) U test (a.k.a. Wilcoxon rank-sum test or Mann-Whitney-Wilcoxon test) is the de facto standard for testing whether two groups are drawn from the same distribution [27, 28]. It is widely applied in medical research [29], for example, when analyzing a biomarker between non-healthy and healthy patients in clinical trials. It is well known (see e.g. [30]) that MW U test is valid for this question, that is, the probability of falsely rejecting the null hypothesis of the two groups being drawn from the same distribution is at most the significance level determined a priori.
@@ 2 Objectives
In this study, we empirically evaluate the validity and power of the MW U test applied to DP-synthetic data in terms of Type I and Type II errors, respectively. On one hand, a test is valid if, for any significance level, the probability that it falsely rejects a correct null hypothesis is no larger than the significance level [31]. If the test is not valid, its use can lead to false scientific discoveries, and hence its practical utility can be even negative. On the other hand, the test’s power refers to the probability of correctly rejecting a false null hypothesis.
In our experiments, we test five different DP-synthetic data generation methods on two real-world medical datasets, as well as data drawn from Gaussian distributions. Our study contributes to understanding the reliability of statistical analysis when DP-synthetic data is utilized as a proxy for private data whose public release is challenging or even impossible.
@@ 3 Methods
In this section, we first present the formal definition of DP. Next, we describe the five DP-synthetic data generation methods used in this study. Finally, we explain the Mann-Whitney U test and how its validity and power can be measured in terms of Type I and Type II errors.
@@ 3.2 DP Methods for Synthetic Data Generation
In this study, we focus on five well-known DP methods for generating synthetic data. These methods have established algorithms or available packages, making them accessible to any practitioner. Following, we provide a brief description of each of these DP methods.
@@
3.3 Mann Whitney U test
The MW U test is a statistical test first proposed by Frank Wilcoxon [40] in 1945 and later, in 1947, formalized by Henry Mann and Donald Whitney [41]. While there are many different uses and interpretations of the test (see e.g. [30] for a comprehensive review), in this paper we focus on the null hypothesis that two samples or groups are drawn from the same distribution. The test carried out on two groups produces a value of the MW U statistic and the corresponding p-value. The U statistic measures the difference between the groups as the number of times an observed member of the first group is smaller than that of the second group, ties being counted as a half time. The p-value indicates the strength of evidence the value of the U statistic provides against the null hypothesis, given that the assumption of the data being independently drawn holds. Following [31], we say that p-value is valid, if it is at most the probability of observing as extreme difference between the groups as the one obtained, for any distribution of the data under the null hypothesis. Consequently, the significance test is valid if its p-value is valid, indicating that the probability of falsely rejecting the null hypothesis is at most the significance level.
The utility of the MW U test on DP-synthetic data can be evaluated in terms of Type I and Type II errors. Let us recall that Type I is the error incurred when a “True” null hypothesis is rejected, producing false inference. On the other hand, Type II is the error of failing to reject a “False” null hypothesis (see Figure 2). A priori selected significance level $\alpha$ defines a threshold that, for any valid hypothesis test, forms an upper bound on the probability of committing Type I error. A typical choice for $\alpha$ is 0.05, indicating a maximum 5% chance of incorrectly rejecting a true null hypothesis. The probability of making a Type II error is often denoted as $\beta$ (beta), from which the power of the test can be determined by computing $1-\beta$. The power of a test can be interpreted as the probability of correctly rejecting a null hypothesis when it is in fact ”False”. The power depends on the analysis task, being affected by factors such as chosen significance level, sample size, and the effect size.
@@ 4 Experimental Evaluation
To empirically evaluate the utility of the MW U test on synthetic data generated by various DP generation methods, we performed a set of experiments using DP-synthetic datasets generated from both real-world datasets and data drawn from Gaussian distributions. As depicted by real distribution in Figure 2, we considered two cases for Gaussian data: one where both groups are drawn from the same distribution (i.e., the null hypothesis is true) and one where they are drawn from distributions with different means (i.e., the null hypothesis is false). In the overall study design (see Figure 1), the real-world and Gaussian datasets correspond to the sensitive data given as input to a DP-synthesizer method that produces a DP-synthetic dataset as output. In the following subsections, we present the datasets, the implementation details of the DP-synthetic data generation methods used, and the experiments conducted.
@@ 4.1 Original Datasets
All the original datasets used in our experiments consist of two variables (i.e., binary- and continuous variable). The binary variable represents the label of two groups (e.g., healthy and non-healthy), while the continuous variable is the one used to compare the groups with the MW U test.
@@ 4.2 Implementations
In our experiments, for all generated DP-synthetic data, the “two-sided” MW U test from SciPy (v1.6.3) package was used to compute their U statistic and p-value. As a point of reference, we also computed the DP-MW U statistic and p-value on the corresponding original sensitive dataset. The DP-MW U test was implemented using Python v3.7 and following the algorithms presented in [19], where 65% of $\epsilon$ and $\delta=10^{-6}$ are used for estimating the size of the smallest group, and the U statistic is privatized using the estimated size and the remaining $\epsilon$.
In the case of the DP Perturbed Histogram, Python v3.7 was also used in the implementation. The noise, added to the original histogram, was sampled from a discrete Laplacian distribution [48] scaled by $\mathrm{\frac{2}{\epsilon}}$, then the noisy counts were normalized by the original dataset size. After that, the synthetic data was obtained by transforming the histogram counts to values using the bin center point. For Private-PGM [49] and MWEM [50], their corresponding open-source packages were used to generate DP-synthetic data. The Private-PGM synthetic data was generated by following the demonstration in Python code presented by McKenna et al. [26] using Laplace distribution scaled by $\frac{2}{split(\epsilon)}$ where $split(\epsilon)$ is the privacy budget ($\epsilon$) divided by the number of marginal queries selected. MWEM was run with default hyperparameters; only $\epsilon$ was changed to show the effect of different privacy budgets. The resulting DP-synthetic data was sampled using the histogram noisy weights returned by the MWEM algorithm. The implementation of DP Smoothed Histogram was also coded in Python v.3.7 following Algorithm 1 provided in the supplement material.
@@ 4.3 Experimental Setup
In the experiments, we investigated the utility of the MW U test at different levels of privacy $\epsilon$. For the DP-MW U test and all DP-synthetic data generation methods, except for DP GAN, we used $\epsilon$ values of 0.01, 0.1, 1, 5, and 10. For the DP GAN experiments, the $\epsilon$ values were 1, 2, 3, 4, 5, and 10. The higher minimum of $\epsilon=1$ was set due to differences between the DP GAN and the other methods. Every experiment was repeated 1000 times, and the proportion of Type I and Type II errors were computed and evaluated at a significance level $\alpha=0.05$.
@@
In all experiments with the GAN discriminator networks a subsampling rate $\gamma$, of 1/500 was used, resulting in mutually exclusive subsets of size 40. The sample size for the GAN training data was 20 000 in all settings and 1000 different generators were trained with models saved at the chosen values of $\epsilon$ (1, 2,3, 4, 5, and 10). Five synthetic datasets of sizes 50, 100, 500, and 1000 were sampled from each of the generators and MW-U tests were conducted on each of these synthetic datasets separately. The DP-hyperparameters were all set to C = 1 for the gradient clipping bound and 1.07 for the noise multiplier as in [25].
@@ 5.2 Real-world Data
Figure 7(a) shows the results of experiments conducted with the prostate cancer dataset. The DP-MW U test performs as expected for an original dataset size of 500 with a group ratio of approx. 50%. The effect of the signal is detected for $\epsilon\geq 1$, while for $\epsilon<1$ the effect is often not detected. Similar behavior is present in DP-synthetic data from DP Perturbed Histogram and MWEM, yet the chance of detecting the effect when $\epsilon<1$ is higher than in the DP-MW U test. In DP-synthetic data from Private-PGM, the signal effect is detected for $\epsilon\geq 5$, with a lower chance of detection for $\epsilon<5$.

### v2

@@ Abstract
Objectives: We investigate how trustworthy are group differences discovered by independent sample tests from DP-synthetic data. The evaluation is carried out in terms of the tests’ Type I and Type II errors. With the former, we can quantify the tests’ validity i.e. whether the probability of false discoveries is indeed below the significance level, and the latter indicates the tests’ power in making real discoveries.
Methods: We evaluate the Mann-Whitney U test, Student’s t-test, chi-squared test and median test on DP-synthetic data. The private synthetic datasets are generated from real-world data, including a prostate cancer dataset (n=500) and a cardiovascular dataset (n=70 000), as well as on bivariate and multivariate simulated data. Five different DP-synthetic data generation methods are evaluated, including two basic DP histogram release methods and MWEM, Private-PGM, and DP GAN algorithms.
Conclusion: A large portion of the evaluation results expressed dramatically inflated Type I errors, especially at privacy budget levels of $\epsilon\leq 1$. This result calls for caution when releasing and analyzing DP-synthetic data: low p-values may be obtained in statistical tests simply as a byproduct of the noise added to protect privacy. A DP smoothed histogram-based synthetic data generation method was shown to produce valid Type I error for all privacy levels tested but required a large original dataset size and a modest privacy budget ($\epsilon\geq 5$) in order to have reasonable Type II error levels.
@@ 1 Introduction
Hypothesis tests for assessing whether two distributions share a certain property are essential tools in analyzing biomedical data. In this work we particularly focus on the Mann-Whitney (MW) U test (a.k.a. Wilcoxon rank-sum test or Mann-Whitney-Wilcoxon test), as it is the de facto standard for testing whether two groups are drawn from the same distribution [27, 28]. It is widely applied in medical research [29], particularly when analyzing a biomarker between non-healthy and healthy patients in clinical trials. It is well known (see e.g. [30]) that MW U test is valid for this question, that is, the probability of falsely rejecting the null hypothesis of the two groups being drawn from the same distribution is at most the significance level determined a priori. Alongside the MW U test, we also consider the Student’s t-test [31], median test [32], and chi-squared test [33]. In general, the choice of statistical test should be guided by the distribution characteristics of the dataset and type of variable under analysis.
@@ 2 Objectives
In this study, we empirically evaluate the validity and power of independent sample tests, such as the MW U test, applied to DP-synthetic data. The Type I and Type II errors are used to measure the test validity and power, respectively. On one hand, a test is valid if, for any significance level, the probability that it falsely rejects a correct null hypothesis is no larger than the significance level [34]. If the test is not valid, its use can lead to false scientific discoveries, and hence its practical utility can be even negative. On the other hand, the test’s power refers to the probability of correctly rejecting a false null hypothesis.
In our experiments with the MW U test, we evaluated five different DP-synthetic data generation methods on bivariate real-world medical datasets, as well as data drawn from two Gaussian distributions. Additionally, we performed experiments with simulated multivariate data to explore the behavior of MW U test, Student’s t-test, median test, and chi-squared test on higher-dimensional DP-synthetic data consisting of different variable types. Our study contributes to understanding the reliability of statistical analysis when DP-synthetic data is utilized as a proxy for private data whose public release is challenging or even impossible.
@@ 3 Methods
In this section, we first present the formal definition of DP. Next, we introduce DP methods for synthetic data generation while describing the five DP-synthesizers used in this study. Following that, we explain the validity and power of a statistical test. Finally, we introduce the independent sample tests considered in this study.
@@ 3.2 DP Methods for Synthetic Data Generation
Generative methods based on marginals share a three-step process: initially, a set of marginals is identified, either manually by a domain expert or through DP-automatic selection. Next, these chosen marginals are measured using DP. Finally, synthetic data is generated from the noisy marginals. To address the challenges of high-dimensional domains, recent methods have been developed to automatically and privately select a subset of marginals ensuring their preservation in the synthetic data generated, such as PrivMRF [41], Privbayes [42], MWEM [22] and AIM [43].
PrivMRF employs Markov Random Fields to generate synthetic data under differential privacy, emphasizing the retention of statistical correlations between selected marginals within the privacy constraints. PrivBayes constructs a Bayesian network under differential privacy, utilizing a selected set of marginals to approximate the underlying data distribution for synthetic data generation. The MWEM algorithm is designed to generate a data distribution that yields query responses closely resembling those obtained from the actual dataset. AIM on the other hand is a workload-adaptive algorithm, allowing for the input of a pre-defined set of marginals to be specifically preserved in the final DP approximated distribution.
There are two approaches to consider when designing a DP workflow: global DP and local DP [13]. Global DP involves aggregating data in a central location and is managed by a trusted curator, ensuring privacy at the dataset level. In contrast, Local DP decentralizes the privacy mechanism by applying it directly to the individual’s data before it is shared. Many applications, such as crowdsourced systems, involve data distributed across multiple individuals who do not trust any other party. These individuals are only willing to share their information if it has been privatized on their own devices prior to transmission. In such cases, local privacy methods such as LoPub and LoCop [44, 45, 46] become applicable, ensuring that each individual’s data remains confidential even when aggregated from diverse sources.
In this study, we focus on five well-known DP methods for generating synthetic data in a global DP setting. These methods have established algorithms or available packages, making them accessible to any practitioner. Following, we provide a brief description of each of these DP methods.
@@
3.3 Validity and Power of Independent Sample Tests
Samples are considered independent when individuals in one group do not influence or share information with individuals in another group. Each group consists of unique members, and no pairing or matching occurs between them. To evaluate potential statistical differences between the two groups, researchers commonly use statistical tests designed for independent samples. These tests determine whether the samples were drawn independently from distributions with shared properties. The independent sample tests considered in this work are the MW U test, Student’s t-test, median test, and chi-squared test.
The validity and power of a statistical test can be evaluated in terms of Type I and Type II errors. Let us recall that Type I is the error incurred when a “True” null hypothesis is rejected, producing false inference. On the other hand, Type II is the error of failing to reject a “False” null hypothesis (see Figure 2). Following [34], we say that p-value corresponding to the observed test statistic is valid, if it is at most the probability of observing as extreme test statistic under the null hypothesis. Consequently, the significance test is valid if its p-value is valid.
A priori selected significance level $\alpha$ defines a threshold that, for any valid hypothesis test, forms an upper bound on the probability of committing Type I error. A typical choice for $\alpha$ is 0.05, indicating a maximum 5% chance of incorrectly rejecting a true null hypothesis. The probability of making a Type II error is often denoted as $\beta$ (beta), from which the power of the test can be determined by computing $1-\beta$. The power of a test can be interpreted as the probability of correctly rejecting a null hypothesis when it is in fact ”False”. The power depends on the analysis task, being affected by factors such as chosen significance level, the effect size, the sample size, and the relative sizes of the different groups. In our experiments, we observed the imbalance between group sizes to have a dominant effect on tests’ power in practice, because of the DP-synthetic data generators’ tendency to produce imbalanced samples for small $\epsilon$ values.
3.4 Mann-Whitney U Test
The MW U test is a statistical test first proposed by Frank Wilcoxon [49] in 1945 and later, in 1947, formalized by Henry Mann and Donald Whitney [50]. While there are many different uses and interpretations of the test (see e.g. [30] for a comprehensive review), in this paper we focus on the null hypothesis that two samples or groups are drawn from the same distribution. The test carried out on two groups produces a value of the MW U statistic and the corresponding p-value. The U statistic measures the difference between the groups as the number of times an observed member of the first group is smaller than that of the second group, ties being counted as a half time. The p-value indicates the strength of evidence the value of the U statistic provides against the null hypothesis, given that the assumption of the data being independently drawn holds.
@@ 3.4 Mann-Whitney U Test
3.5 Student’s t-Test, Chi-squared Test and Median Test
The Student’s t-test (independent or unpaired t-test) [31] is a widely utilized parametric statistical test that assesses whether the means of two independent samples are significantly different. The null hypothesis states that the means are statistically equivalent, while the alternative hypothesis suggests that they are not. The test is valid for two independent samples if their distributions are normal and their variances are equal.
The chi-squared test [33], is a non-parametric test used to analyze the association of two categorical variables by utilizing a contingency table. Under the null hypothesis, the observed (joint) frequencies should equal to expected (marginal) frequencies, meaning that the variables are independent. Since, under the null hypothesis, the test statistic approximately follows a Chi-squared distribution, the test’s validity depends on the sample size. However, for small $n$ and for $2\times 2$ tables, the appropriate alternative is the Fisher’s exact test.
Median test [32] is a nonparametric method used to test the null hypothesis of two (or more) independent samples being drawn from distributions of equal medians. The test is valid as long as the distributions have equal densities in the neighborhood of the median (see e.g. [51] and references therein).
@@ 4 Experimental Evaluation
To empirically evaluate the utility of independent sample tests applied to DP-synthetic datasets, we conducted a set of experiments. In each experiment, either simulated or real-world data was used to represent the original sensitive dataset. This data was subsequently used to train DP-synthetic data generation methods. Finally, the independent sample tests were carried out on synthetic data produced by the generator.
First, we examined the behavior of MW U test on DP-synthetic data generated based on bivariate real-world datasets or simulated data drawn from Gaussian distributions. As depicted by real distribution in Figure 2, we considered two cases for Gaussian data: one where both groups are drawn from the same distribution (i.e., the null hypothesis is true) and one where they are drawn from distributions with different means (i.e., the null hypothesis is false). While in practice synthesizing datasets consisting of only two variables would have quite limited use cases, these experiments allow demonstrating the fundamentals of how different DP synthetization approaches affect the validity and power of statistical tests. In order to provide a more realistic setup, we further performed experiments on a simulated multivariate dataset. The validity and power of the MW U test, Student’s t-test, median test, and chi-squared test were explored in these experiments.
In the overall study design (see Figure 1), the real-world, Gaussian, and simulated multivariate datasets correspond to the sensitive data given as input to a DP-synthesizer method that produces a DP-synthetic dataset as output. In the following subsections, we present the datasets, the implementation details of the DP-synthetic data generation methods used, and the experiments conducted.
@@ 4.1 Original Datasets
First, we experimented with a setup, where the sensitive original data consists of only two variables (i.e., a binary variable and a continuous variable). The binary variable represents group membership (e.g. healthy or non-healthy), while the continuous variable is the one used to compare the groups with the MW U test.
@@
Finally, we experimented with simulated multivariate datasets. The simulation was based on the real-world prostate cancer dataset. The included variables were the patient’s age, PSA level, prostate volume, the use of 5-alpha-reductase inhibitors (5-ARIs) medication, prostate imaging reporting and data systems (PIRADS) score, and a class label indicating low-risk or high-risk PCa. The simulated datasets were generated by a GaussianCopulaSynthesizer from the Synthetic Data Vault (SDV) [58] trained on the real-world dataset. In the SVD settings, the age variable was configured to follow a normal distribution, while the remaining numerical variables were set to follow a beta distribution. In experiments with a false null hypothesis, SVD was conditioned to generate simulated datasets with an equal number of high-risk and low-risk patients. For experiments with a true null hypothesis, the condition was to generate only one class (low-risk) for the simulated dataset, and subsequently, half of the data were randomly assigned to the high-risk class.
@@ 4.2 Implementations
In our experiments, for the generated DP-synthetic data, we used the hypothesis tests provided by the Scipy v1.6.3 package [59], such as the mannwhitneyu function for the MW U test. We used two-sided tests, with all the tests statistics and p-values computed using the Scipy function’s default values. As a point of reference, we also computed the DP-MW U statistic and p-value on the corresponding original sensitive dataset. The DP-MW U test was implemented using Python v3.7 and following the algorithms presented in [19], where 65% of $\epsilon$ and $\delta=10^{-6}$ are used for estimating the size of the smallest group, and the U statistic is privatized using the estimated size and the remaining $\epsilon$.
In the case of the DP Perturbed Histogram, Python v3.7 was also used in the implementation. The noise, added to the original histogram, was sampled from a discrete Laplacian distribution [60] scaled by $\mathrm{\frac{2}{\epsilon}}$, then the noisy counts were normalized by the original dataset size. After that, the synthetic data was obtained by transforming the histogram counts to values using the bin center point. For Private-PGM [61] and MWEM [62], their corresponding open-source packages were used to generate DP-synthetic data. The Private-PGM synthetic data was generated by following the demonstration in Python code presented by McKenna et al. [26] using Laplace distribution scaled by $\frac{2}{split(\epsilon)}$ where $split(\epsilon)$ is the privacy budget ($\epsilon$) divided by the number of marginal queries selected. MWEM was run with default hyperparameters; only $\epsilon$ was changed to show the effect of different privacy budgets. The resulting DP-synthetic data was sampled using the histogram noisy weights returned by the MWEM algorithm. The implementation of DP Smoothed Histogram was also coded in Python v.3.7 following Algorithm 1 provided in the supplement material. In all our experiments, the DP-synthetic data generators were configured to preserve all the one-way and two-way marginals.
@@ 4.3 Experimental Setup
In the experiments, we investigated the utility of the statistical test at different levels of privacy $\epsilon$. For the DP-MW U test and all DP-synthetic data generation methods, except for DP GAN, we used $\epsilon$ values of 0.01, 0.1, 1, 5, and 10. For the DP GAN experiments, the $\epsilon$ values were 1, 2, 3, 4, 5, and 10. The higher minimum of $\epsilon=1$ was set due to differences between the DP GAN and the other methods. Every experiment was repeated 1000 times, and the proportion of Type I and Type II errors were computed and evaluated at a significance level $\alpha=0.05$.
@@
In all experiments with the GAN discriminator networks a subsampling rate $\gamma$, of 1/500 was used, resulting in mutually exclusive subsets of size 40. The sample size for the GAN training data was 20 000 in all settings and 1000 different generators were trained with models saved at the chosen values of $\epsilon$ (1, 2,3, 4, 5, and 10). Five synthetic datasets of sizes 50, 100, 500, and 1000 were sampled from each of the generators and MW U tests were conducted on each of these synthetic datasets separately. The DP-hyperparameters were all set to C = 1 for the gradient clipping bound and 1.07 for the noise multiplier as in [25].
@@
Setup for Simulated Multivariate Data
In the experiments with a simulated multivariate dataset, we considered the Private-PGM and MWEM synthesizers. Using the generated DP-synthetic data, we empirically assessed the proportion of Type I and Type II errors for the MW U test for an ordinal variable (PI-RADS score), Student’s t-test for a normally distributed continuous variable (age), median test for another continuous variable (PSA), and chi-squared test of independence for a binary variable (use of 5ARIs medication).
For these experiments, we generated 1000 simulated multivariate datasets for dataset sizes of 50, 100, 500, 1000, and 20000. Subsequently, for each simulated dataset, we applied the DP-synthetic data generator at each epsilon value. The proportion of Type I and Type II error were measured across the DP-synthetic datasets, with the condition that the requirements for running the statistical test were met in at least 50 of the generated DP-synthetic datasets (see supplementary material A.2 for further details on cases when tests are undefined, such as when the DP-synthetic data consists of only single class).
@@ 5.2 Real-world Data
Figure 7(a) shows the results of experiments conducted with the prostate cancer dataset. The DP-MW U test performs as expected for an original dataset size of 500 with a group ratio of approx. 50%. The null hypothesis is rejected for $\epsilon\geq 1$, while for $\epsilon<1$ it is often not rejected. Similar behavior is present in DP-synthetic data from DP Perturbed Histogram and MWEM, yet the chance of rejecting the null hypothesis when $\epsilon<1$ is higher than in the DP-MW U test. In DP-synthetic data from Private-PGM, the null hypothesis is rejected for $\epsilon\geq 5$ more often than for $\epsilon<5$.
@@
5.3 Simulated Multivariate Data
Figure 9: Proportion of Type I error conditioned to the number of DP-synthetic datasets where the statistical test is feasible; a) MW U test applied to an ordinal variable (PI-RADS score); b) Student’s T-test on a normally distributed variable (Age); c) Median test on continuous variable (PSA); d) Chi-squared test on a binary variable (5ARI medication).
Figure 10: Proportion of Type II error conditioned to the number of DP-synthetic datasets where the statistical test is feasible; a) MW U test applied to an ordinal variable (PI-RADS score); b) Student’s T-test on a normally distributed variable (Age); c) Median test on continuous variable (PSA); d) Chi-squared test on a binary variable (5ARI medication).
In Figure 9, the proportion of Type I errors for various statistical tests (i.e., MW U test, Student’s t-test, median test, and chi-squared test) is presented. From these results, we observe that false discoveries are also prone to occur similarly to the previous experiments with only two variables. The validity of the tests is preserved only for largest tested privacy budgets combined with large amounts of the original sensitive data. Same kind of trend was observed for all statistical tests under consideration. For Private-PGM, a substantial drop in Type I error was observed for $\epsilon=0.01$ and dataset size $<20000$. On closer examination, we observed that with the smallest privacy budgets the size of the smaller of the two groups tends to be very small or even zero. This can be seen from the numbers of times the test requirements failed, as presented in supplementary material A.2, where the tests fail when the size of smaller group is zero. The power of all evaluated tests strongly depend on the group size imbalance in the sample, so that for a fixed sample size they have the highest power for equal group sizes and the power shrinks to zero when the smaller group size goes to zero. Therefore, the tendency of the low privacy budgets to produce imbalanced samples counters the tendency to produce fake group differences to some extent.
In the case of Type II error proportions (Figure 10), the results depend on the magnitude of the group differences in the original data, how it is preserved by the GaussianCopulaSynthesizer, and the size of the simulated dataset. As a baseline or point of reference, we first present the Type II error probabilities computed over the 1000 simulated multivariate datasets that represent the original sensitive data before the synthetic data is generated based on them. Then, we illustrate the corresponding Type II error probabilities for the DP-synthetic data with different privacy budgets. For the synthetic data, especially for $\epsilon<10$, we observe that the Type II errors are often lower than those of on the original data, indicating that true group differences are discovered more often from the synthetic data than from the original. However, this is explained perfectly by the large Type I error probabilities presented in Figure 9, indicating that the fake group differences present in the synthetic data are so strong that they end up getting discovered rather than the true ones that are too weak to be discovered from the original data. For $\epsilon=0.01$, the large Type II error of Private-PGM also mirrors the low Type I error, caused by the loss of power due to the group size imbalance.